OpenStudy (anonymous):

how do I find y-intercept given two points my two points are (0,1) and (4,0) and my slope is -4 I think

OpenStudy (anonymous):

The standard form for the equation of a line is y=mx+b, where m is your slope and b is your y-intercept. You know that the line passes through (0,1) and (4,0). First, you need to calculate your slope. m = (y2-y1) / (x2-x1) = (0-1) / (4-0) = -1/4 So we can plug slope into our equation: y = (-1/4)x + b Now, you can pick either of your points and plug them into your equation and solve for b.